OpenStudy (anonymous):

csc=5/3 and tan= 3/4. find cos, sec, and cot

OpenStudy (anonymous):

cossec = 5/3 = 1/sin --> sin(x) = 3/5 sin^2 + cos^2 = 1 --> cos^2 = 25/25 - 9/25 = 16/25 --> cos(x) = 4/5 sec = 1/cos -> sec = 5/4 cot = 1/tg --> cot = 4/3

OpenStudy (anonymous):

cot x = 1/tanx so cot x = 4/3 if tan = 3/4 then cos = 4/5 sec = 1/cos = 5/4

OpenStudy (anonymous):

cos x = 4/5 because of 5,4,3 sided right angled triangle

OpenStudy (anonymous):

how to find tan 53 or sin 47 or cos 47??????